Your answer is partially correct. Try again. At the end of 2016, Metlock, Inc. has accounts receivable of $666,300 and an allowance for doubtful accounts of $24,740. On January 24, 2017, it is learned that the company’s receivable from Madonna Inc. is not collectible and therefore management authorizes a write-off of $4,398. (a) Prepare the journal entry to record the write-off. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit Entry field with correct answer Allowance for Doubtful Accounts Entry field with correct answer 4,398 Entry field with correct answer Entry field with correct answer Accounts Receivable Entry field with correct answer Entry field with correct answer 4,398 (b) What is the cash realizable value of the accounts receivable before the write-off and after the write-off? Before Write-Off After Write-Off Cash realizable value $Entry field with incorrect answer 641550 $Entry field with incorrect answer
Answer #1

Journal entry

Date account and explanation debit credit
Allowance for doubtful accounts 4398
Account receivable 4398
(To record write off)

(b) What is the cash realizable value of the accounts receivable before the write-off and after the write-off?

Before write off After write off
Cash realizable value 666300-24740 = 641560 (666300-4398)-(24740-4398) = 641560